OpenStudy (anonymous):

s=1/2n(a+K) could someone show me how to do this? im confused because i thought you had to multiply the entire equation by 2 but thats giving me wrong answer

OpenStudy (freckles):

There isn't enough information in this question.

OpenStudy (anonymous):

solving for K

OpenStudy (freckles):

\[s=\frac{1}{2}n(a+K) ?\] If this is right, yes multiplying 2 as a first step is good. But that still will not isolate K yet. You will still have to perform more steps. Is this what you get when you multiply 2 on both sides? 2s=n(a+K) try isolating (a+k) next by undoing the multiplication by n to it.

OpenStudy (anonymous):

Yeah I get that part I'm just wondering why I dont multiply a and K by 2 and it ends up being 2s=n(a+K) instead

OpenStudy (freckles):

are you asking why 2/2=1?

OpenStudy (anonymous):